OpenStudy (anonymous):

Which of the following expressions is equal to (x+2)+(x-2)(2x+1)? A.2x^2-2x B.2x^2-4x C.2x^2+x D.4x^2+2x

OpenStudy (anonymous):

Easiest way to check: Plug in x=1 for x in the first equation, then do the same for the other four. If you get the same answer then that's it.

OpenStudy (anonymous):

ya i plugged in x=1 and i got 3 for all of them but what would be the answer....?

OpenStudy (anonymous):

Or choose another number, something obscure, like 12. But you can also find it by expanding out and combining like terms. (x+2) + (x-2)(2x+1) = (x+2) + (2x^2 + x -4x -2) = 2x^2 -2x A.

OpenStudy (anonymous):

ok thankx!!!